OpenStudy (anonymous):

use a half-angle formula to find the exact value of each expression sin15

OpenStudy (aum):

\[ \sin^2(\frac a2) = \frac 12\{1 - \cos(a)\} \]

OpenStudy (aum):

Set a = 30 degrees in the above half-angle formula and compute sin(15).

OpenStudy (anonymous):

cos 30 \[\frac{ \sqrt{3} }{ 2}\]

OpenStudy (aum):

correct. \[ \sin^2(15) = \frac 12\{1 - \frac{\sqrt 3}{2}\} = \frac 14\{2-\sqrt {3}\} \\ \sin(15) = \frac 12\sqrt{2- \sqrt{3}} \]

OpenStudy (anonymous):

im confused can i show you my work?

OpenStudy (aum):

Go ahead.

OpenStudy (anonymous):